OpenStudy (anonymous):

is the following equation true? and plz when solving explain including the properties used

OpenStudy (anonymous):

first of all it is not an equation...... equation means LHS = RHS here there is no equal sign and no RHS

OpenStudy (anonymous):

3 log 2 x +\[3 \log _{2} x +1/2 \log _{2} y – 3 \log _{2} z = \log 2 \sqrt[x^3]{y}\div z^{3}\] log 2 y – 3 log 2 z = log 2 .

OpenStudy (anonymous):

the attachment was the wrong one